The ratio of Julio's money to Marcus's money was 7:10. After Julio bought 4 books at $4.50 each the ratio of Julio's money to Marcus's money was 1:4. Use a bar graph to find how much money Julioo and Marcus have left.

Answer by stanbon(75887)

The ratio of Julio's money to Marcus's money was 7:10. After Julio bought 4 books at $4.50 each the ratio of Julio's money to Marcus's money was 1:4. Use a bar graph to find how much money Julioo and Marcus have left.
----
Original amounts of money
Julio has 7x dollars
Marcus has 10x dollars
------
Julio spent 4*4.50 = 18 dollars on books.
---
Equation:
(7x-18)/10x = 1/4
---
28x - 72 = 10x
18x = 72
x = 4
---
Julio has 7x-18 = 28-18 = 10 dollars
Marcus has 10x = 40 dollars
======================================
Comment: Bar graphs cannot be displayed on this site.
======================================
Cheers,
Stan H.